# RDA-INR: Riemannian Diffeomorphic Autoencoding via Implicit Neural Representations

### v2

@@ 3.2 Diffeomorphic registration and the LDDMM Riemannian distance
In this work, we focus on the LDDMM diffeomorphic registration framework. The objective of LDDMM is to find a diffeomorphism $\phi:\Omega\rightarrow\Omega$ that matches two objects $O_{1}$ and $O_{2}$ (e.g., images or shapes defined on the spatial domain $\Omega\subset\mathbb{R}^{d}$) by deforming $O_{1}$ into $O_{2}$. More precisely, LDDMM wants $\phi\cdot O_{1}=O_{2}$ with $\phi\cdot O$ a left group action of the diffeomorphism group on the set of objects. For instance, an image $I:\Omega\rightarrow\mathbb{R}^{n}$ can be deformed by a left group action $\phi\cdot I:=I\circ\phi^{-1}$ and a shape $\mathcal{S}:=\{x\mid x\text{ on the shape}\}$ can be deformed by a left group action $\phi\cdot\mathcal{S}:=\phi(\mathcal{S})$.
To construct a diffeomorphism that achieves $\phi\cdot O_{1}=O_{2}$, LDDMM considers a subgroup of the diffeomorphism group. This subgroup consists of diffeomorphisms emerging as flows of ordinary differential equations (ODEs). Define a time-dependent vector field $v:\Omega\times[0,1]\rightarrow\mathbb{R}^{d}$ and assume $v(\cdot,t)\in V$ for some Banach space $V$. The corresponding ODE is $\frac{\mathrm{d}x}{\mathrm{d}t}=v(x,t)$ and it gives us a flow of diffeomorphisms:
$\frac{\partial\phi_{t}}{\partial t}(x)$ $=v(\phi_{t}(x),t),$ (3)
$\phi_{0}(x)$ $=x.$
To make these ODE flows diffeomorphisms, one has to enforce smoothness on the velocity vector fields $v(\cdot,t)\in V$. Generally, this is accomplished by choosing an appropriate Banach space $V$ and assuming $v\in L^{1}([0,1],V)$. One class of such spaces are admissible Banach spaces:
Definition 1 (Admissible Banach spaces [30]).
Let $C_{0}^{1}(\Omega,\mathbb{R}^{d})$ be the Banach space of continuously differentiable vector fields $\nu$ on the open and bounded domain $\Omega\subset\mathbb{R}^{d}$ such that both $\nu$ and its Jacobian $J\nu$ vanish on $\partial\Omega$ and at infinity. Furthermore, define a Banach space $V\subset C_{0}^{1}(\Omega,\mathbb{R}^{d})$ and let $\lVert\nu\rVert_{1,\infty}:=\lVert\nu\rVert_{\infty}+\lVert J\nu\rVert_{\infty}$ for $\nu\in V$. The Banach space $V$ is admissible if it is (canonically) embedded in $(C_{0}^{1}(\Omega,\mathbb{R}^{d}),\lVert\cdot\rVert_{1,\infty})$. In other words, there exist a constant $C$ such that $\forall\nu\in V$, $\lVert\nu\rVert_{V}\geq C\lVert\nu\rVert_{1,\infty}$.
@@
$d_{G}(\phi,\psi):=\inf_{v\in L^{1}([0,1],V)}\quad$ $\left(\int_{0}^{1}\lVert v(\cdot,t)\rVert_{V}\mathrm{d}t\right)$
@@
where $\lVert\cdot\rVert_{V}$ is a norm on $V$. The LDDMM approach selects the $\phi\in G$ such that $\phi\cdot O_{1}=O_{2}$ and $d_{G}(\text{id},\phi)$ is small. Alternatively, in case the objects $O_{1}$ and $O_{2}$ belong to a set $\mathcal{O}$ of diffeomorphic objects, we can reformulate this problem as finding a distance between the two objects $O_{1}$ and $O_{2}$:
Theorem 1 ([30]).
[…]
@@
$d_{\mathcal{O}}(O_{1},O_{2})=\inf_{v(\cdot,t)}\quad$ $\left(\int_{0}^{1}\lVert v(\cdot,t)\rVert_{V}\mathrm{d}t\right)$ (5)
@@
[…]
One particular example when $\phi\rightarrow\phi\cdot O_{\text{temp}}$ is continuous as a mapping from $G$ to $\mathcal{O}$ is when considering the shape $O_{\text{temp}}=\{x\mid f(x)=0)\}=\partial A$ for some continuous function $f$ and $A\subset\mathbb{R}^{d}$ some non-empty compact set. In this case $\phi\cdot O_{\text{temp}}:=\phi(O_{\text{temp}})$ and we use the topology induced by the Hausdorff distance on $\mathcal{O}$. For more in-depth information on the case of shapes seen as submanifolds of $\mathbb{R}^{d}$, we refer to Bauer et al. [77]. Additionally, when considering images as functions $I:\Omega\rightarrow\mathbb{R}^{n}$ and using the group action $\phi\cdot I:=I\circ\phi^{-1}$, $d_{\mathcal{O}}$ is a distance for a specific class of $\lVert\cdot\rVert_{V}$ [26, 78].
As mentioned above, the goal of LDDMM is to approximately solve Equations (4) and (5). However, it is known that the solution $\phi_{t}$ to Equation (5) is reparameterization invariant. Hence, there exists an infinite number of $t\rightarrow\phi_{t}$ that trace out the same curve in the diffeomorphism group and that all have the same integral cost. To resolve this issue, LDDMM minimizes an energy that is not reparameterization invariant:
Theorem 2 ([30]).
[…]
$E_{G}(\phi,\psi):=\inf_{v\in L^{1}([0,1],V)}\quad$ $\left(\int_{0}^{1}\lVert v(\cdot,t)\rVert_{V}^{2}\mathrm{d}t\right)$ (6)
@@
$\psi=\phi_{1}\circ\phi.$
Then $E_{G}(\phi,\psi)=d_{G}(\phi,\psi)^{2}$. Moreover, we define the induced energy $E_{\mathcal{O}}$ on $\mathcal{O}$ as
[…]
or alternatively:
$E_{\mathcal{O}}(O_{1},O_{2})=\inf_{v(\cdot,t)}\quad$ $\left(\int_{0}^{1}\lVert v(\cdot,t)\rVert_{V}^{2}\mathrm{d}t\right)$ (7)
$\phi_{1}\cdot O_{1}=O_{2}.$
As $E_{G}(\phi,\psi)=d_{G}(\phi,\psi)^{2}$, $E_{\mathcal{O}}(O_{1},O_{2})=d_{\mathcal{O}}(O_{1},O_{2})^{2}$. More precisely, it can be shown that any $\phi_{t}$ solving the energy minimization problem (7) solves the distance problem in Equation (5). Conversely, any constant speed curve $t\rightarrow\phi_{t}$ (i.e., $\lVert v(\cdot,t)\rVert_{V}$ stays constant over time) that solves the distance problem in Equation (5) solves the energy minimization problem in Equation (7).
Using Theorem 2, LDDMM solves optimization problem (7) instead of optimization problem (5). To solve the energy minimization problem (7), LDDMM solves the relaxed problem:
$\inf_{v(\cdot,t)}\quad$ $\mathcal{D}(\phi_{1}\cdot O_{1},O_{2})+\sigma^{2}\int_{0}^{1}\lVert v(\cdot,t)\rVert_{V}^{2}\mathrm{d}t$
where $\mathcal{D}$ is some data fidelity term and $\sigma\in\mathbb{R}$. For instance, when dealing with images, one can use $\mathcal{D}(I_{1},I_{2})=\lVert I_{1}-I_{2}\rVert_{L_{2}(\Omega)}^{2}$ and $\phi\cdot I=I\circ\phi^{-1}$.
@@ 3.3 Diffeomorphic latent modeling via LDDMM PGA
Besides registering two objects $O_{1}$ and $O_{2}$, the Riemannian distance in Theorem 1 can be used for statistical latent modeling. Specifically, LDDMM is combined with principal geodesic analysis (PGA) for diffeomorphic latent modeling. In the original papers, a Bayesian approach introduces LDDMM PGA. However, we introduce LDDMM PGA from a slightly different perspective, namely closer to the original PGA formulation in [41].
The basic idea behind PGA is to first estimate a Fréchet mean and subsequently estimate a geodesic submanifold that best explains the data. For finding the Fréchet mean within the LDDMM PGA framework, we consider the set $\mathcal{O}:=\{\phi\cdot O_{\text{temp}}\mid\phi\in G\}$ for some reference template $O_{\text{temp}}$. This reference template is arbitrary as for every $O\in\mathcal{O}$, we have $\mathcal{O}=\{\phi\cdot O\mid\phi\in G\}$, which means that any $O$ can be seen as the template. In LDDMM PGA, the Fréchet mean is a template $\mathcal{T}\in\mathcal{O}$ that minimizes the variance of the data. More precisely, the Fréchet mean should solve:
$\min_{\mathcal{T}}\quad$ $\operatorname*{\mathop{\mathbb{E}}}_{O\sim\rho}\left[d_{\mathcal{O}}(\mathcal{T},O)^{2}\right],$
with $\rho$ the distribution of objects $O$ and $d_{\mathcal{O}}$ given as the distance in Equations (4) and (5). Using Theorem 2, we can also minimize:
$\min_{\mathcal{T}}\quad$ $\operatorname*{\mathop{\mathbb{E}}}_{O\sim\rho}\left[E_{\mathcal{O}}(\mathcal{T},O)\right].$
Writing out $E_{\mathcal{O}}$ using Equation (7) and approximating the expectation using $N$ available samples $\{O_{i}\}_{i=1}^{N}$, we get:
$\min_{\{v_{i}(\cdot,t)\}_{i=1}^{N},\mathcal{T}}\quad$ $\frac{1}{N}\sum_{i=1}^{N}\left[\int_{0}^{1}\lVert v_{i}(\cdot,t)\rVert_{V}^{2}\mathrm{d}t\right]$ (8)
@@
An optimization over a time-dependent velocity field is difficult. To circumvent this issue, the diffeomorphism can be parameterized via an initial velocity vector field $v_{0}$. More precisely, if $V$ is a Hilbert space, the extrema of Equations (6) and (7) satisfy the so-called EPDiff equation. This equation maps an initial velocity vector field $v_{0}\in V$ to a time-dependent velocity vector field $v(\cdot,t)$ that satisfies $\int_{0}^{1}\lVert v(\cdot,t)\rVert_{V}^{2}\mathrm{d}t=\lVert v_{0}(\cdot)\rVert_{V}^{2}$. Hence, defining $\phi_{t}$ as the flow of diffeomorphisms corresponding to $v(\cdot,t)$, we can construct a map $\text{exp}:V\rightarrow G$ as $\text{exp}{(v_{0})}:=\phi_{1}$. The exp map allows optimization over initial velocity vector fields and it has been used for, among others, images [79, 35, 36, 31, 80] and landmarks [28, 34]. While providing an in-depth treatment of the EPDiff equation is outside the scope of this work, for our discussion here, it is sufficient to know that the initial velocities constitute the tangent space at the identity diffeomorphism and that there exists an exponential map $\text{exp}{(v_{0})}$ that maps an initial velocity vector field to a diffeomorphism. This situation is similar to PGA where an initial tangent vector is mapped to an element on the manifold via the exponential map. For more information on the EPDiff equation, we refer to Younes [30, 81].
Using the exp map, we replace the ODE in Equation (8) with $\phi^{i}:=\text{exp}{(v_{i,0})}$ for some initial velocity vector field $v_{i,0}\in V$. In addition, using the properties of the exp map, we replace $\int_{0}^{1}\lVert v_{i}(\cdot,t)\rVert_{V}^{2}\mathrm{d}t$ by $\lVert v_{i,0}\rVert_{V}^{2}$. Finally, we replace the hard constraint $\phi_{1}^{i}\cdot\mathcal{T}=O_{i}$ by a soft penalty $\mathcal{D}(\phi^{i}\cdot\mathcal{T},O_{i})$. Combining these adjustments yields:
$\min_{\{v_{i,0}\}_{i=1}^{N},\mathcal{T}}\quad$ $\frac{1}{N}\sum_{i=1}^{N}\left[\mathcal{D}(\phi^{i}\cdot\mathcal{T},O_{i})+\sigma^{2}\lVert v_{i,0}\rVert_{V}^{2}\right]$ (9)
@@
where $\sigma\in\mathbb{R}$. To optimize the initial velocity vector fields and the template, one creates discrete initial velocities $v_{i,0}$ and a discrete version of the template. For instance, the template can be represented by a finite-resolution image or as a mesh.
Now that we know how to solve for the Fréchet mean in LDDMM PGA, we only need to find a geodesic submanifold that best explains the data. LDDMM PGA restricts the discretized initial velocities to a linear subspace via a, possibly orthonormal, matrix $W$ and a diagonal matrix $\Lambda$. Putting this into Equation (9) yields the LDDMM PGA problem:
$\min_{W,\Lambda,\{z_{i}\}_{i=1}^{N},\mathcal{T}}\quad$ $\frac{1}{N}\sum_{i=1}^{N}\left[\mathcal{D}(\phi^{i}\cdot\mathcal{T},O_{i})+\sigma^{2}\lVert v_{i,0}\rVert_{V}^{2}\right]$ (10)
@@
Additional loss components related to priors on $W$, $\Lambda$, and the latent vectors can be added to the above optimization problem. These priors relate to the standard Bayesian derivation to obtain the LDDMM PGA model [35, 36]. However, these considerations are outside the scope of this work.

### v3

@@ 3.2 Diffeomorphic registration and the LDDMM Riemannian distance
In this work, we focus on the Large Deformation Diffeomorphic Metric Mapping (LDDMM) registration framework. The objective of LDDMM is to find a diffeomorphism $\phi\colon\Omega\rightarrow\Omega$ that matches two objects $O_{1}$ and $O_{2}$ (e.g., images or shapes defined on the spatial domain $\Omega\subset\mathbb{R}^{d}$) by deforming $O_{1}$ into $O_{2}$. More precisely, LDDMM wants $\phi\cdot O_{1}=O_{2}$ with $\phi\cdot O$ a left group action of the diffeomorphism group on the set of objects. This group action formulation is very flexible due to the ambient diffeomorphisms. For instance, a function $I\colon\Omega\rightarrow\mathbb{R}^{n}$ representing an image or an implicit shape representation can be deformed by a left group action $\phi\cdot I:=I\circ\phi^{-1}$. Moreover, a set-based definition of a shape $\mathcal{S}:=\{x\mid x\text{ on the shape or pointcloud}\}$ can be deformed by $\phi\cdot\mathcal{S}:=\phi(\mathcal{S})$.
To construct a diffeomorphism achieving $\phi\cdot O_{1}=O_{2}$, LDDMM considers a subgroup of the diffeomorphism group. This subgroup consists of diffeomorphisms emerging as flows of ordinary differential equations (ODEs). Define a time-dependent vector field $v\colon\Omega\times[0,1]\rightarrow\mathbb{R}^{d}$ and assume $v(\cdot,t)\in V$ for some Banach space $V$ with norm $\left\lVert\cdot\right\rVert_{V}$. The corresponding ODE is $\frac{\mathrm{d}x}{\mathrm{d}t}=v(x,t)$ and gives us a flow $\phi_{t}$:
(3) $\frac{\partial\phi_{t}}{\partial t}(x)$ $=v(\phi_{t}(x),t),\quad\phi_{0}(x)$ $=x.$
To make these ODE flows diffeomorphisms, one has to enforce smoothness on the velocity vector fields $v(\cdot,t)\in V$. Generally, this is accomplished by choosing an appropriate Banach space $V$ and assuming $v\in L^{1}([0,1],V)$. One class of such spaces is admissible Banach spaces:
Definition 3.1 (Admissible Banach spaces [80]).
Let $C_{0}^{1}(\Omega,\mathbb{R}^{d})$ be the Banach space of continuously differentiable vector fields $\nu$ on the open and bounded domain $\Omega\subset\mathbb{R}^{d}$ such that both $\nu$ and its Jacobian $J\nu$ vanish on $\partial\Omega$. Furthermore, define a Banach space $V\subset C_{0}^{1}(\Omega,\mathbb{R}^{d})$ and let $\left\lVert\nu\right\rVert_{1,\infty}:=\left\lVert\nu\right\rVert_{\infty}+\left\lVert J\nu\right\rVert_{\infty}$ for $\nu\in C_{0}^{1}(\Omega,\mathbb{R}^{d})$. The Banach space $V$ is admissible if it is (canonically) embedded in $(C_{0}^{1}(\Omega,\mathbb{R}^{d}),\left\lVert\cdot\right\rVert_{1,\infty})$. In other words, there exist a constant $C$ such that $\forall\nu\in V$, $\left\lVert\nu\right\rVert_{V}\geq C\left\lVert\nu\right\rVert_{1,\infty}$.
@@
$d_{G}(\phi,\psi)^{2}:=\inf_{v\in L^{2}([0,1],V)}$ $\left(\int_{0}^{1}\left\lVert v(\cdot,t)\right\rVert_{V}^{2}\mathrm{d}t\right)$
@@
where $\left\lVert\cdot\right\rVert_{V}$ is a norm on $V$. The LDDMM approach selects the $\phi\in G$ such that $\phi\cdot O_{1}=O_{2}$ and $d_{G}(\text{id},\phi)$ is small. Alternatively, when $O_{1}$ and $O_{2}$ belong to a set $\mathcal{O}$ of diffeomorphic objects, we can reformulate this problem as finding a distance between $O_{1}$ and $O_{2}$:
Theorem 3.2 ([80]).
Assume that $V$ is an admissible Banach space. Then $(G,d_{G})$ is a complete metric space. Furthermore, assume we consider a set $\mathcal{O}:=\{\phi\cdot O_{\text{temp}}\mid\phi\in G\}$ for some template object $O_{\text{temp}}$. Then we can define a pseudo-distance $d_{\mathcal{O}}(O_{1},O_{2})$ on $\mathcal{O}$ as
(4) $d_{\mathcal{O}}(O_{1},O_{2}):=\inf_{\phi}\left(d_{G}({\text{id}},\phi)\mid O_{2}=\phi\cdot O_{1},\phi\in G\right)$
@@
(5) $d_{\mathcal{O}}(O_{1},O_{2})^{2}=\inf_{v(\cdot,t)}$ $\left(\int_{0}^{1}\left\lVert v(\cdot,t)\right\rVert_{V}^{2}\mathrm{d}t\right)$
@@
In addition, $d_{\mathcal{O}}$ is a distance if the action $\phi\rightarrow\phi\cdot O_{\text{temp}}$ is continuous from $G$ to $\mathcal{O}$ given the topology induced by $d_{G}$ on $G$ and some topology on $\mathcal{O}$.
One example when $\phi\rightarrow\phi\cdot O_{\text{temp}}$ is continuous as a mapping from $G$ to $\mathcal{O}$ is when considering the shape $O_{\text{temp}}=\{x\mid f(x)=0\}=\partial A$ for some continuous function $f$ and $A\subset\mathbb{R}^{d}$ a non-empty compact set. In this case $\phi\cdot O_{\text{temp}}:=\phi(O_{\text{temp}})$ and we use the topology induced by the Hausdorff distance on $\mathcal{O}$. For more in-depth information on the case of shapes seen as submanifolds of $\mathbb{R}^{d}$, we refer to Bauer et al. [9]. Moreover, for functions $I\colon\Omega\rightarrow\mathbb{R}^{n}$ and $\phi\cdot I:=I\circ\phi^{-1}$, $d_{\mathcal{O}}$ is a distance for a specific class of $\left\lVert\cdot\right\rVert_{V}$ [11, 49].
As mentioned above, the goal of LDDMM is to approximately solve Equations (4) and (5). As the hard constraint in Equation (5) is difficult to enforce, LDDMM introduces a soft penalty penalizing the difference between $\phi_{1}\cdot O_{1}$ and $O_{2}$. In addition, this soft penalty allows us to consider objects that are not diffeomorphic to each other. The soft penalty yields the relaxed problem:
$\inf_{v(\cdot,t)}$ $\mathcal{D}(\phi_{1}\cdot O_{1},O_{2})+\sigma^{2}\int_{0}^{1}\left\lVert v(\cdot,t)\right\rVert_{V}^{2}\mathrm{d}t$
@@
where $\mathcal{D}$ is some data fidelity term and $\sigma\in\mathbb{R}$. For instance, when dealing with images, one can use $\mathcal{D}(I_{1},I_{2})=\left\lVert I_{1}-I_{2}\right\rVert_{L_{2}(\Omega)}^{2}$ and $\phi\cdot I=I\circ\phi^{-1}$, while for meshes we can use a varifold data attachment term [18] with $\phi\cdot\mathcal{S}:=\phi(\mathcal{S})$ and $\mathcal{S}:=\{x\mid x\text{ on the shape}\}$. We note that the term $\mathcal{D}$ is very flexible. For instance, in Section 4.1, we discuss a data attachment term $\mathcal{D}(O_{1},O_{2})$ for the case where $O_{1}$ is represented by an implicit representation and $O_{2}$ by an oriented point cloud.
@@ 3.3 Diffeomorphic latent modeling via LDDMM PGA
Besides registering two objects $O_{1}$ and $O_{2}$, the Riemannian distance in Theorem 3.2 can be used for statistical latent modeling. Specifically, LDDMM is combined with principal geodesic analysis (PGA) for diffeomorphic latent modeling. In the original papers, a Bayesian approach introduces LDDMM PGA [81, 84]. However, we introduce LDDMM PGA from a slightly different perspective, namely closer to the original PGA formulation in [29].
The basic idea behind PGA is to first estimate a Fréchet mean and subsequently estimate a geodesic submanifold that best explains the data. For finding the Fréchet mean within the LDDMM PGA framework, we consider the set $\mathcal{O}:=\{\phi\cdot O_{\text{temp}}\mid\phi\in G\}$ for some reference template $O_{\text{temp}}$. This reference template is arbitrary as for every $O\in\mathcal{O}$, we have $\mathcal{O}=\{\phi\cdot O\mid\phi\in G\}$, which means that any $O$ can be seen as the template. In LDDMM PGA, the Fréchet mean is a template $\mathcal{T}\in\mathcal{O}$ that minimizes the variance of the data:
$\min_{\mathcal{T}}\operatorname*{\mathop{\mathbb{E}}}_{O\sim\rho}\left[d_{\mathcal{O}}(\mathcal{T},O)^{2}\right]$
with $\rho$ the distribution of objects $O$ and $d_{\mathcal{O}}$ given as the distance in Equations (4) and (5). Writing out $d_{\mathcal{O}}$ using Equation (5) and approximating the expectation using $N$ available samples $\{O_{i}\}_{i=1}^{N}$, we get:
(6) $\min_{\{v_{i}(\cdot,t)\}_{i=1}^{N},\mathcal{T}}$ $\frac{1}{N}\sum_{i=1}^{N}\left[\int_{0}^{1}\left\lVert v_{i}(\cdot,t)\right\rVert_{V}^{2}\mathrm{d}t\right]$
@@
An optimization over a time-dependent velocity field is difficult. To circumvent this issue, we note that the geodesic optimization problems between $O_{i}$ and $\mathcal{T}$, which appear in Equation (6), are equivalent to an optimization problem over initial velocity vector fields $v_{i0}$. This equivalence is due to the diffeomorphism subgroup $G$ having a Lie group structure with Lie Algebra $V$. More precisely, if $V$ is a Hilbert space, the geodesics in $G$, which solve optimization problems (4) and (5), are characterized by a Riemannian exponential map. This exponential map is characterized by the so-called EPDiff equations and maps an initial vector field $v_{0}\in V$ to a time-dependent vector field $v(\cdot,t)$ that satisfies $\int_{0}^{1}\left\lVert v(\cdot,t)\right\rVert_{V}^{2}\mathrm{d}t=\left\lVert v_{0}(\cdot)\right\rVert_{V}^{2}$. Hence, defining $\phi_{t}$ as the flow of diffeomorphisms corresponding to $v(\cdot,t)$, we can construct a map $\text{exp}\colon V\rightarrow G$ as $\text{exp}{(v_{0})}:=\phi_{1}$. Providing an in-depth treatment of the EPDiff equation is outside the scope of this work. For more information on the EPDiff equation, we refer to Younes [80, 79]. For our discussion here, it is sufficient to note that replacing the time-dependent velocity vector fields by the $\phi^{i}:=\text{exp}{(v_{i,0})}$ and replacing $\int_{0}^{1}\left\lVert v_{i}(\cdot,t)\right\rVert_{V}^{2}\mathrm{d}t$ by $\left\lVert v_{i,0}\right\rVert_{V}^{2}$ yields an equivalent formulation of optimization problem (6).
We make these adjustments and replace the hard constraint $\phi_{1}^{i}\cdot\mathcal{T}=O_{i}$ by a soft penalty $\mathcal{D}(\phi^{i}\cdot\mathcal{T},O_{i})$:
(7) $\min_{\{v_{i,0}\}_{i=1}^{N},\mathcal{T}}$ $\frac{1}{N}\sum_{i=1}^{N}\left[\mathcal{D}(\phi^{i}\cdot\mathcal{T},O_{i})+\sigma^{2}\left\lVert v_{i,0}\right\rVert_{V}^{2}\right]$
@@
where $\sigma\in\mathbb{R}$. LDDMM PGA [81, 84] considers image data and optimizes the initial velocity vector fields and the template by using a grid-based discretization for $v_{i,0}$ and a finite-resolution image for the template. The exponential map on the discretized initial velocities is obtained by solving a discretized EPDiff equation. While to the best of our knowledge LDDMM PGA has not explored mesh data, Equation (7) allows a mesh template and spatially discretized $v_{i,0}$ via kernel-based Reproducing Kernel Hilbert Spaces (RKHS). This approach results in an exponential map calculation that preserves the discretization [18, 47, 80].
Now that we know how to solve for the Fréchet mean in LDDMM PGA, we only need to find a geodesic submanifold that best explains the data. While we earlier mentioned that PGA [29] estimates the Fréchet mean first and then finds a geodesic submanifold minimizing the variance, LDDMM PGA estimates them jointly. LDDMM PGA restricts the discretized initial velocities to a linear subspace via a, possibly orthonormal, matrix $W$ and a diagonal matrix $\Lambda$. Putting this into Equation (7) yields the LDDMM PGA problem:
(8) $\min_{W,\Lambda,\{z_{i}\}_{i=1}^{N},\mathcal{T}}$ $\frac{1}{N}\sum_{i=1}^{N}\left[\mathcal{D}(\phi^{i}\cdot\mathcal{T},O_{i})+\sigma^{2}\left\lVert v_{i,0}\right\rVert_{V}^{2}\right]$
@@
Additional loss terms related to priors on $W$, $\Lambda$, and $z_{i}$ can be added to optimization problem (8). These priors relate to the standard Bayesian derivation of the LDDMM PGA model [81, 84]. However, these considerations are outside the scope of this work.